OpenStudy (atomicreaper):

Find the solution of this system of equations. separate the x-and y-values with a comma. x=8+y x-11y=-12

OpenStudy (welshfella):

from equation 2 x = 11y - 12 but x = 8 + y also so by the transitive property of equality 11y - 12 = 8 + y 11y - y = 8 + 12 10y = 20 y = 2 Now plug y = 2 into the equation x = 8 + y to find the value of x

OpenStudy (atomicreaper):

Ok thank you

OpenStudy (welshfella):

yw